# suppSupplementary References

### v3

@@ Abstract
Previous works on Human Pose and Shape Estimation (HPSE) from RGB images can be broadly categorized into two main groups: parametric and non-parametric approaches. Parametric techniques leverage a low-dimensional statistical body model for realistic results, whereas recent non-parametric methods achieve higher precision by directly regressing the 3D coordinates of the human body mesh. This work introduces a novel paradigm to address the HPSE problem, involving a low-dimensional discrete latent representation of the human mesh and framing HPSE as a classification task. Instead of predicting body model parameters or 3D vertex coordinates, we focus on predicting the proposed discrete latent representation, which can be decoded into a registered human mesh. This innovative paradigm offers two key advantages. Firstly, predicting a low-dimensional discrete representation confines our predictions to the space of anthropomorphic poses and shapes even when little training data is available. Secondly, by framing the problem as a classification task, we can harness the discriminative power inherent in neural networks. The proposed model, VQ-HPS, predicts the discrete latent representation of the mesh. The experimental results demonstrate that VQ-HPS outperforms the current state-of-the-art non-parametric approaches while yielding results as realistic as those produced by parametric methods when trained with little data. VQ-HPS also shows promising results when training on large-scale datasets, highlighting the significant potential of the classification approach for HPSE. See the project page at https://g-fiche.github.io/research-pages/vqhps/
Figure 1: VQ-HPS formulates the human pose and shape estimation problem as a classification task in a vector-quantized latent space. We present the results of VQ-HPS on two challenging scenarios with in-the-wild conditions and poor illumination, comparing its performance to that of HMR [1], Cliff [2] and FastMETRO [3] when trained on little data.
@@ 1 Introduction
Capturing and understanding human motion from RGB data is a fundamental task in computer vision, with many applications such as character animation for the movie and video-game industries [4, 5, 6] or performance optimization in sports [7, 8]. However, due to depth ambiguity, estimating 3D human pose and shape from monocular images is an underdetermined problem. To overcome this issue, parametric approaches (also called model-based) use statistical models of the human body, which enable the reconstruction of a human 3D mesh by predicting a small number of parameters [9, 10, 11, 12, 13]. Earlier methods were optimization-based, estimating the parameters of a human body model iteratively using 2D cues [14, 15, 16]. However, their need for a good initialization, slow running time, and propensity to converge towards local minima led many recent works to focus on regression-based methods, which predict the parameters of a human body model directly from RGB data [1, 2, 17]. Despite producing realistic results in most scenarios, methods regressing the parameters of a human body model face several issues well documented in the literature: 1) Parametric methods struggle in capturing detailed body shape and are biased towards the mean shape [18]; 2) Most human body models use rotations along the kinematic tree for expressing the pose. In addition to being difficult to predict for neural networks [19, 20], this representation induces error accumulation when all rotations are predicted simultaneously [21, 22]; 3) Most regression methods extract global feature vectors from the image as an input, which do not contain fine-grained local details [23].
@@
This work introduces a method significantly different from all prior human pose and shape estimation (HPSE) approaches. Instead of predicting the parameters of a human body model or 3D coordinates, we learn to predict a discrete latent representation of 3D meshes, transforming the HPSE into a classification problem in which we can exploit the originally targeted discriminative power of Transformers, which has been proven unmatched in natural language processing. For learning our discrete latent representation of meshes, we build on the vector quantized-variational autoencoder (VQ-VAE) [27] framework and adapt it to the fully convolutional mesh autoencoder proposed in [28]. The encoder of the proposed model, called Mesh-VQ-VAE, provides a low-dimensional discrete latent representation preserving the spatial structure of the mesh. We then propose a Transformer-based encoder-decoder model, called VQ-HPS, for learning to solve the HPSE problem using the cross-entropy loss. Once the mesh discrete representation is predicted, we can decode it using the pre-trained Mesh-VQ-VAE decoder and obtain a full mesh following the SMPL mesh topology [9]. Since the Mesh-VQ-VAE is pre-trained on a large human motion capture database [29], it automatically learns to decode smooth and realistic human meshes. This is particularly interesting when training with little data: VQ-HPS learns to predict sequences of indices corresponding to realistic meshes early in the training process as demonstrated in Appendix C.
In the context of few training data availability, VQ-HPS achieves state-of-the-art performance on the challenging in-the-wild 3DPW [30] and EMDB [31] benchmarks: it significantly outperforms other methods quantitatively while producing qualitative results as realistic as parametric methods (see Fig. 1). Moreover, it shows promising results when trained on standard large-scale datasets, enhancing the significant potential of the classification-based approach for solving the HPSE problem.
@@
  • Code and trained models will be released upon acceptance.
@@ 4.1 Proposed HPSE method
Mesh-VQ-VAE. For learning discrete representations of meshes, we build on the fully convolutional mesh autoencoder [28] (see Sec. 3) for encoding the full canonical mesh vertices $V_{c}\in\mathbb{R}^{6890\times 3}$ to a latent representation $z\in\mathbb{R}^{N\times L}$. We add a vector quantization step in the latent space similar to [27] (see Sec. 3), which maps $z$ to the discrete latent representation $z_{q}\in\{1,\dots,S\}^{N}$. While the fully convolutional architecture preserves the spatial structure of the mesh, the added quantization step allows us to view the HPSE as a classification task as we aim to predict the indices of the latent mesh representation given an image. Our Mesh-VQ-VAE Fig. 2 can be seen as a VQ-VAE [27] whose architecture corresponds to the fully convolutional mesh autoencoder.
@@
Feature extractors. The first step for image-based HPSE is to extract features from the image. We use CNN backbones to preserve the spatial structure of the image, and we obtain features $X\in\mathbb{R}^{H\times W\times C}$, where $C$ is the number of channels of the backbone and $H$ and $W$ are the spatial dimension sizes. We use two feature extractors. The feature extractor of the camera and rotation predictors gives $X_{r}ot$. The feature extractor gives $X_{mesh}$ with $W=1$ in the latent canonical mesh predictor.
@@
Rotation and camera prediction. We start by predicting the mesh rotation and the perspective camera parameters (see again Fig. 2). These predictions depend on the image features and an initial body pose $p\in\mathbb{R}^{17\times 3}$ following the Human3.6M [77] joints layout. We predict the rotation $\hat{R}$ and the weak perspective camera parameters $\hat{\pi}$.
@@
Latent canonical mesh regressor encoder. We then predict the discrete latent representation of the canonical mesh. The Transformer encoder inputs are the features extracted by the CNN backbone. Before being fed to the Transformer encoder, we apply a 1x1 convolution on the image features to make them of dimension $X^{\prime}_{mesh}\in\mathbb{R}^{H\times W\times D}$ where $D$ is the hidden state size of the Transformer. These features are flattened to obtain $HW$ tokens of dimension $D$, and then we add positional encoding. The obtained tokens are fed to a Transformer encoder, using self-attention between all image tokens to output encoded image features $X_{feat}\in\mathbb{R}^{HW\times D}$.
@@
Latent canonical mesh regressor decoder. The Transformer decoder takes as inputs $N$ learned mesh tokens $M_{t}$ of size $D$, each responsible for predicting an index of the Mesh-VQ-VAE discrete latent representation. We need to solve $N$ classification problems, one for each index. Each problem is a $S$-class classification, $S$ corresponding to the size of the Mesh-VQ-VAE codebook. The Transformer decoder consists of self-attention between learned tokens and cross-attention with image features. It outputs latent mesh features $z_{feat}\in\mathbb{R}^{N\times D}$. Then (see Fig. 2), to obtain the logits $\hat{z}_{log}\in\mathbb{R}^{N\times S}$, we rely on the mesh features as well as on the previously predicted rotation and camera. We obtain the predicted discrete representation $\hat{z}_{q}\in\{1,\dots,S\}^{N}$ by applying an $\arg\max(\cdot)$ operation.
@@ 4.2 Training VQ-HPS
where $\hat{s}$ is the predicted scale, $\hat{t}$ the predicted 2D translation and $\hat{J}_{3D}$ are the 3D joints computed from the predicted oriented mesh vertices $\hat{\mathcal{V}}$. $\Pi$ is the orthographic projection using the matrix $\begin{bmatrix}1&0&0\\ 0&1&0\end{bmatrix}^{\top}$ and $J_{2D}$ are the ground truth 2D joints.
@@ 5.3 Training on limited data
Figure 4: Qualitative results We compare our method with HMR [1], Cliff [2] and FastMETRO [3] on 3DPW trained on 3DPW (first 3 rows), and EMDB trained on EMDB.
We train VQ-HPS separately on the 3DPW, COCO, and EMDB training sets (see Sec. 5.1) to see how it performs when trained on limited data. We compare our performance with HMR [1], Cliff [2], and FastMETRO [3] trained with the same data. We chose these 3 models for comparison because HMR is the basic architecture for parametric human mesh recovery, Cliff is the state-of-the-art (SOTA) for parametric HPSE, and FastMETRO is the SOTA for non-parametric HPSE and the closest method to ours. For these experiments, the backbone for all networks is ResNet-50 [80] pre-trained on ImageNet [81]. We use the public implementation of FastMETRO and adapt the HMR and Cliff implementations provided by [60]. For all comparisons, we use FastMETRO-S, as this version is the closest to VQ-HPS. When training on COCO, we propose an improved version of VQ-HPS, replacing the MLP of the latent canonical mesh regressor with a Transformer implementing self-attention between the latent mesh features $z_{feat}$. Quantitative results are shown in Tab. 1, and visualizations are available in Fig. 4.
Overall, VQ-HPS outperforms the SOTA methods significantly when training on 3DPW and EMDB (see Tab. 1, col. 1 and 3). Visualization of the results confirms that our method performs best (see Fig. 4), and we propose a more detailed analysis of the error in Appendix D. HMR and Cliff show realistic predictions but are less accurate than VQ-HPS. Despite rather good metrics, FastMETRO produces non-smooth results that do not correspond to human body shapes. This is probably due to the limited training sets, as the results displayed in the original paper looked more realistic. This highlights a clear limitation of non-parametric approaches predicting the 3D coordinates; in comparison, VQ-HPS needs much less data to provide realistic predictions, probably because learning sequences of indices corresponding to anthropomorphic meshes is easier than understanding the 3D structure of the 3D vertices. Furthermore, VQ-HPS benefits from the large-scale pre-training on human motion datasets, acting as a regularization to allow learning with fewer image data labeled with 3D poses.
@@ 5.4 Training on large-scale datasets
Table 2: Comparison to the SOTA methods. We evaluate VQ-HPS trained on a mix of datasets (see Sec. 5.1) with standard metrics on 3DPW and EMDB and compare them to the state of the art. On the left part, methods use a ResNet-50 backbone. On the right, models use an HRNet backbone, except TokenHMR [76] that uses a ViT [49] backbone and additional data. All results are given in mm.
@@ Acknowledgments
This study is part of the EUR DIGISPORT project supported by the ANR within the framework of the PIA France 2030 (ANR-18-EURE-0022). This work was performed using HPC resources from the “Mésocentre” computing center of CentraleSupélec, École Normale Supérieure Paris-Saclay, and Université Paris-Saclay supported by CNRS and Région Île-de-France.
@@ Appendix A Implementation details
Appendix B Qualitative results
Some qualitative results obtained with VQ-HPS trained on real image datasets and finetuned on the 3DPW training set (see Sec. 5.4, quantitative results in the "With finetuning" part of Sec. 5.4) are shown in Fig. 1.

### v4

@@ Abstract
Previous works on Human Pose and Shape Estimation (HPSE) from RGB images can be broadly categorized into two main groups: parametric and non-parametric approaches. Parametric techniques leverage a low-dimensional statistical body model for realistic results, whereas recent non-parametric methods achieve higher precision by directly regressing the 3D coordinates of the human body mesh. This work introduces a novel paradigm to address the HPSE problem, involving a low-dimensional discrete latent representation of the human mesh and framing HPSE as a classification task. Instead of predicting body model parameters or 3D vertex coordinates, we focus on predicting the proposed discrete latent representation, which can be decoded into a registered human mesh. This innovative paradigm offers two key advantages. Firstly, predicting a low-dimensional discrete representation confines our predictions to the space of anthropomorphic poses and shapes even when little training data is available. Secondly, by framing the problem as a classification task, we can harness the discriminative power inherent in neural networks. The proposed model, VQ-HPS, predicts the discrete latent representation of the mesh. The experimental results demonstrate that VQ-HPS outperforms the current state-of-the-art non-parametric approaches while yielding results as realistic as those produced by parametric methods when trained with few data. VQ-HPS also shows promising results when training on large-scale datasets, highlighting the significant potential of the classification approach for HPSE. See the project page at https://g-fiche.github.io/research-pages/vqhps/.
Figure 1: VQ-HPS formulates the human pose and shape estimation problem as a classification task in a vector-quantized latent space. We present the results of VQ-HPS on two challenging scenarios with in-the-wild conditions and poor illumination, comparing its performance to that of HMR [1], CLIFF [2] and FastMETRO-S [3] when trained on little data.
@@ 1 Introduction
Capturing and understanding human motion from RGB data is a fundamental task in computer vision, with many applications such as character animation for the movie and video-game industries [4, 5, 6] or performance optimization in sports [7, 8]. However, due to depth ambiguity, estimating 3D human pose and shape from monocular images is an underdetermined problem. To overcome this issue, parametric approaches (also called model-based) use statistical models of the human body, which enable the reconstruction of a 3D human mesh by predicting a small number of parameters [9, 10, 11, 12, 13]. Earlier methods were optimization-based, estimating the parameters of a human body model iteratively using 2D cues [14, 15, 16]. However, their need for a good initialization, slow running time, and propensity to converge towards local minima led many recent works to focus on regression-based methods, which predict the parameters of a human body model directly from RGB data [1, 2, 17]. Despite producing realistic results in most scenarios, methods regressing the parameters of a human body model face several issues well documented in the literature: 1) Parametric methods struggle in capturing detailed body shape and are biased towards the mean shape [18]; 2) Most human body models use rotations along the kinematic tree for expressing the pose. In addition to being difficult to predict for neural networks [19, 20], this representation induces error accumulation when all rotations are predicted simultaneously [21, 22]; 3) Most regression methods extract global feature vectors from the image as an input, which do not contain fine-grained local details [23].
@@
This work introduces a method significantly different from all prior human pose and shape estimation (HPSE) approaches. Instead of predicting the parameters of a human body model or 3D coordinates, we learn to predict a discrete latent representation of 3D meshes, transforming the HPSE into a classification problem in which we can exploit the originally targeted discriminative power of Transformers, which has been proven unmatched in natural language processing. For learning our discrete latent representation of meshes, we build on the vector quantized-variational autoencoder (VQ-VAE) [27] framework and adapt it to the fully convolutional mesh autoencoder proposed in [28]. The encoder of the proposed model, called Mesh-VQ-VAE, provides a low-dimensional discrete latent representation preserving the spatial structure of the mesh. We then propose a Transformer-based encoder-decoder model, called VQ-HPS, for learning to solve the HPSE problem using the cross-entropy loss. Once the mesh discrete representation is predicted, we can decode it using the pre-trained Mesh-VQ-VAE decoder and obtain a full mesh following the SMPL mesh topology [9]. Since the Mesh-VQ-VAE is pre-trained on a large human motion capture database [29], it automatically learns to decode smooth and realistic human meshes. This is particularly interesting when training with little data: VQ-HPS learns to predict sequences of indices corresponding to realistic meshes early in the training process, as demonstrated in the supplementary materials.
In the context of few training data availability, VQ-HPS achieves state-of-the-art performance on the challenging in-the-wild 3DPW [30] and EMDB [31] benchmarks: it significantly outperforms other methods quantitatively while producing qualitative results as realistic as parametric methods (see Fig. 1). Moreover, it also shows SOTA results when trained on standard large-scale datasets, enhancing the significant potential of the classification-based approach for solving the HPSE problem.
@@
  • Code and trained models are available from the project page.
@@ 4.1 Proposed HPSE method
Mesh-VQ-VAE. For learning discrete representations of meshes, we build on the fully convolutional mesh autoencoder [28] (see Sec. 3) for encoding the full canonical mesh vertices $V_{c}\in\mathbb{R}^{6890\times 3}$ to a latent representation $z\in\mathbb{R}^{N\times L}$. We add a vector quantization step in the latent space similar to [27] (see Sec. 3), which maps $z$ to the discrete latent representation $z_{q}\in\{1,\dots,S\}^{N}$. While the fully convolutional architecture preserves the spatial structure of the mesh, the added quantization step allows us to view the HPSE as a classification task as we aim to predict the indices of the latent mesh representation given an image. Our Mesh-VQ-VAE in Fig. 2 can be seen as a VQ-VAE [27] whose architecture corresponds to the fully convolutional mesh autoencoder.
@@
Feature extractors. The first step for image-based HPSE is to extract features from the image. We use CNN backbones to preserve the spatial structure of the image, and we obtain features $X\in\mathbb{R}^{H\times W\times C}$, where $C$ is the number of channels of the backbone and $H$ and $W$ are the spatial dimension. We use two feature extractors. The feature extractor of the camera and rotation predictors gives $X_{r}ot$. The feature extractor gives $X_{mesh}$ with $W=1$ in the latent canonical mesh predictor.
@@
Rotation and camera prediction. We start by predicting the mesh rotation and the perspective camera parameters (see again Fig. 2). These predictions depend on the image features and an initial body pose $p\in\mathbb{R}^{17\times 3}$ following the Human3.6M [77] joints layout and corresponding to the SMPL T-pose. We predict the rotation $\hat{R}$ and the weak perspective camera parameters $\hat{\pi}$.
@@
Latent canonical mesh regressor encoder. We then predict the discrete latent representation of the canonical mesh. The Transformer encoder inputs are the features extracted by the CNN backbone. Before being fed to the Transformer encoder, we apply a 1x1 convolution on the image features to make them of dimension and obtain $X^{\prime}_{mesh}\in\mathbb{R}^{H\times W\times D}$ where $D$ is the hidden state size of the Transformer. These features are flattened to obtain $HW$ tokens of dimension $D$, and then we add positional encoding. The obtained tokens are fed to a Transformer encoder, using self-attention between all image tokens to output encoded image features $X_{feat}\in\mathbb{R}^{HW\times D}$.
@@
Latent canonical mesh regressor decoder. The Transformer decoder takes as inputs $N$ learned mesh tokens $M_{t}$ of size $D$, each responsible for predicting an index of the Mesh-VQ-VAE discrete latent representation. We need to solve $N$ classification problems, one for each index. Each problem has $S$ classes, $S$ corresponding to the size of the Mesh-VQ-VAE codebook. The Transformer decoder consists of self-attention between learned tokens and cross-attention with image features. It outputs latent mesh features $z_{feat}\in\mathbb{R}^{N\times D}$. Then (see Fig. 2), to obtain the logits $\hat{z}_{log}\in\mathbb{R}^{N\times S}$, we rely on the mesh features as well as on the previously predicted rotation and camera. We obtain the predicted discrete representation $\hat{z}_{q}\in\{1,\dots,S\}^{N}$ by applying an $\arg\max(\cdot)$ operation.
@@ 4.2 Training VQ-HPS
where $\hat{s}$ is the predicted scale, $\hat{t}$ the predicted 2D translation and $\hat{J}_{3D}$ are the 3D joints computed from the predicted oriented mesh vertices $\hat{V}$. $\Pi$ is the orthographic projection using the matrix $\begin{bmatrix}1&0&0\\ 0&1&0\end{bmatrix}^{\top}$ and $J_{2D}$ denotes the ground truth 2D joints.
@@ 5.3 Training on limited data
Figure 4: Qualitative results We compare our method with HMR [1], CLIFF [2] and FastMETRO-S [3] on 3DPW trained on 3DPW (first 3 rows), and EMDB trained on EMDB.
We train VQ-HPS separately on the 3DPW, COCO, and EMDB training sets (see Sec. 5.1) to see how it performs when trained on limited data. We compare our performance with HMR [1], CLIFF [2], and FastMETRO [3] trained with the same data. We chose these 3 models for comparison because HMR is the basic architecture for parametric human mesh recovery, CLIFF is the SOTA for parametric HPSE, and FastMETRO is the SOTA for non-parametric HPSE and the closest method to ours. For these experiments, the backbone for all networks is ResNet-50 [80] pre-trained on ImageNet [81]. We use the public implementation of FastMETRO and adapt the HMR and CLIFF implementations provided by [60]. For all comparisons, we use FastMETRO-S, as this version is the closest to VQ-HPS. When training on COCO, we propose an improved version of VQ-HPS, replacing the MLP of the latent canonical mesh regressor with a Transformer implementing self-attention between the latent mesh features $z_{feat}$. Quantitative results are shown in Tab. 1, and visualizations are available in Fig. 4.
Overall, VQ-HPS outperforms the SOTA methods significantly when training on 3DPW and EMDB (see Tab. 1, col. 1 and 3). Visualization of the results confirms that our method performs best (see Fig. 4), and we propose a more detailed analysis of the error in Appendix D. HMR and CLIFF show realistic predictions but are less accurate than VQ-HPS. Despite rather good metrics, FastMETRO produces non-smooth results that do not correspond to human body shapes. This is probably due to the limited training sets, as the results displayed in the original paper looked more realistic. This highlights a clear limitation of non-parametric approaches predicting the 3D coordinates; in comparison, VQ-HPS needs much less data to provide realistic predictions, probably because learning sequences of indices corresponding to anthropomorphic meshes is easier than understanding the structure of the 3D vertices. Furthermore, VQ-HPS benefits from the large-scale pre-training on human motion datasets, acting as a regularization to allow learning with fewer image data labeled with 3D poses.
@@ 5.4 Training on large-scale datasets
Table 2: Comparison to the SOTA methods. We evaluate VQ-HPS trained on a mix of datasets without finetuning on 3DPW (see Sec. 5.1) with standard metrics on 3DPW and EMDB and compare them to the state of the art. On the left part, methods use a ResNet-50 backbone. On the right, models use an HRNet backbone, except TokenHMR [76] that uses a ViT [49] backbone and additional data. All results are given in mm.
@@ Acknowledgments
This study is part of the EUR DIGISPORT project supported by the ANR within the framework of the PIA France 2030 (ANR-18-EURE-0022). This work was performed using HPC resources from the “Mésocentre” computing center of CentraleSupélec, École Normale Supérieure Paris-Saclay, and Université Paris-Saclay supported by CNRS and Région Île-de-France. This work has been partially supported by MIAI@Grenoble Alpes, (ANR-19-P3IA-0003).
@@ Appendix A Implementation details
Appendix B Additional qualitative results
Figure 1: Additional comparisons. We compare our method with HMR, CLIFF, and FastMETRO-S on 3DPW trained on 3DPW (first row) and EMDB trained on EMDB.
Additional comparisons with other methods trained on little data are provided in Fig. 1 to complement the results of Fig. 4.
@@ Appendix B Additional qualitative results
We also provide some qualitative results obtained with VQ-HPS using an HRNet backbone as evaluated in Sec. 5.4) are shown in Fig. 2.